# AnTKV: Anchor Token-Aware Sub-Bit Vector Quantization for KV Cache in Large Language Models

### v1

@@ 3.2 Online Anchor Token Selection
In the online stage, the gradient can no longer be used as AnS due to prohibitive memory and computational costs under real-time constraints. To this end, we here give the error propagation analysis of the self-attention (i.e., Attn). Specifically, the error propagation analysis is based on the perturbation bound of Attn that is related to each row of $\bm{K}$ and $\bm{V}$; see Theorem 3.1 (see Appendix B for proof).
@@
We remark that the error propagation factors corresponding to $\bm{K}_{j,:}$ and $\bm{V}_{j,:}$ given in Theorem 3.1 can be regarded as the upper bound of the gradient of the self-attention operator related to $\bm{K}_{j,:}$ and $\bm{V}_{j,:}$.
The computation involving $\bm{K}$ in Theorem 3.1 entails substantial overhead, rendering it impractical for use during online inference. To address this, we propose a simplified variant that excludes the contribution of $\bm{V}$ to the quantization error of $\bm{K}$. By this simplification, we obtain the following reformulation. A detailed evaluation of the effectiveness of AnS is provided in Appendix C.
$\text{AnS}(\bm{K}_{j,:})=\sum_{i}\bm{A}_{i,j}(1-\bm{A}_{i,j})\cdot\left\|\bm{Q}_{i,:}\right\|_{2}$ (5)
During the prefill phase, AnS serves as an effective metric to identify anchor tokens that preserve critical contextual information by measuring their impact on attention outputs. In contrast, although AnS can still be computed during decoding, the anchor token it identifies may have already been quantized during prefill, preventing the preservation of its FP16 value for reducing accumulated errors. An important observation is that $\bm{A}$ and $\bm{A}\odot(\bm{E}-\bm{A})$ exhibit strong locality during decoding (see Appendix D), with anchor tokens primarily concentrated at the beginning and end of the sequence. In particular, the beginning tokens—referred to as sink tokens—are almost always identified as anchor tokens during prefill, indicating that our method effectively handle sink tokens. Motivated by this locality, we adopt a sliding window approximation of AnS during decoding.
@@ 3.3 Implementation
For the offline stage, the gradient of $\bm{K}$ and $\bm{V}$ is used as a weight to learn centroids. We implement it with a custom LinearWithAct to capture KV cache and corresponding gradients. Subsequently, we employ the weighted K-Means provided by cuML to perform efficient clustering.
For the online stage, AnS is derived from the error propagation factor given in Theorem 3.1 and (5). We design and implement a dedicated GPU kernel using Triton to efficiently compute it in conjunction with FlashAttention, enabling long-context inference. Since AnS requires reduction operations over the attention score matrix $\bm{A}$ and its transformed form $\bm{A}\odot(\bm{E}-\bm{A})$ along query (column-wise), directly fusing its computation into FlashAttention is impossible. To preserve the efficiency of FlashAttention, we decouple the AnS computation and execute it immediately afterward. To support this design, we extend FlashAttention to preserve three auxiliary tensors: the $L_{2}$ norm of each query vector, the key-wise (row-wise) sum, and the key-wise maximum of the matrix $\bm{Q}\bm{K}^{T}$. These additional outputs enable efficient reconstruction of the attention scores and allow AnS to be computed with minimal overhead. More details about the implementation are described in Algorithm 1 of Appendix E.
@@ 4 Experiments
In this section, we compare the performance of AnTKV with existing KV quantization methods through extensive experiments. The experimental setup, including models, benchmark datasets, evaluation metrics, and selected baselines, is as follows.
Models, Datasets, Metrics, and Parameter Settings. To validate the effectiveness of AnTKV in KV cache quantization and ensure the generality of results, we evaluate five representative models from the LLaMA and Mistral families: LLaMA-2-7B, LLaMA-3-8B, LLaMA-3-8B-Instruct, Mistral-7B, and Mistral-7B-Instruct-v0.3. We use 128 samples of length 2048 from the WikiText2-v1 training set as the calibration. Perplexity is evaluated on WikiText-2 and C4, while zero-shot accuracy on MMLU hendrycks2021mmlu , ARC-C clark2018arc , MathQA amini2019mathqa , and PIQA bisk2020piqa is used to assess understanding and reasoning. Long-context performance is measured using LongBench bai2024longbench . For perplexity and zero-shot benchmarks, quantized KV cache is employed for attention. In LongBench, the FP16 KV cache is first utilized to compute attention outputs, before quantization is applied.
[…]
@@ 4.1 Perplexity Results
Perplexity is a standard metric that is widely used to evaluate the quality of the output of LLMs, with lower values indicating better performance. The perplexity results for different KV quantization approaches on WikiText-2 and C4 are presented in Table 1. The results in this table indicate that the proposed AnTKV consistently achieves competitive or superior perplexity across various bit-widths and model architectures. Under 4-bit and 2-bit quantization, AnTKV achieves competitive or superior performance compared to existing approaches. In the 1-bit and sub-bit regimes, AnTKV significantly outperforms all baselines. It is worth mentioning that on the C4 dataset under sub-bit quantization, the baseline methods exhibit extremely high perplexity due to anchor tokens not being well captured by fixed centroids. Thanks to effective AnS design and innovative Anchor Token selection, AnTKV substantially reduces perplexity, e.g., from 66.28 to 14.42 (LLaMA-3-8B, 0.75-bit).
Table 1: The evaluations are conducted using each model’s maximum context length, specifically 4096 for LLaMA-2-7B and 8192 for LLaMA-3-8B and Mistral-7B. "Ours" refers to our method (AnTKV) without using any anchor tokens, while Ours-1% denotes AnTKV where 1% of the tokens are treated as anchor tokens and retained in FP16. For simplicity, the reported bit-widths do not consider the effect of centroids.
@@ 4.2 Understanding and Reasoning Benchmark
To ensure that AnTKV exhibits broad understanding and reasoning capabilities, we evaluate its performance on four representative benchmarks: MMLU, ARC-Challenge, PIQA, and MathQA, using both LLaMA-8B-Instruct and Mistral-7B-Instruct. These benchmarks respectively assess multi-domain knowledge reasoning (MMLU), complex question answering (arc_challenge), commonsense reasoning (PIQA), and mathematical problem-solving (MathQA). Figure 4 illustrates the comparison results of AnTKV with SKVQ and CQ. Similarly, it can be seen that AnTKV consistently outperforms SKVQ and CQ across all four benchmarks under low-bit quantization, especially at 1-bit and sub-bit levels.
Figure 4: Understanding and Reasoning ability evaluation accuracy results on MMLU, ARC-C, PIQA, and MathQA under different quantization bit-widths.
@@ 4.3 Long-Context Benchmark
To validate the effectiveness of AnTKV in handling long-context, We conduct several experiments on the LLaMA-8B-Instruct model using the LongBench benchmark, a diverse collection of tasks such as question answering, retrieval, and summarization, designed to systematically evaluate long-context understanding in language models. We report results on eleven representative sub-tasks from LongBench, along with averaged performance. Due to alignment issues of KIVI and SKVQ, we exclude the triviaqa and gov_report sub-tasks from the comparison. As shown in Figure 5, AnTKV preserves nearly FP16 at 4- and 2-bit quantization across almost all tasks. At the 1-bit quantization, the performance of KIVI and SKVQ has a significant drop. In contrast, AnTKV and CQ still maintain a relatively high accuracy. To further investigate the robustness under aggressive compression, we compare AnTKV and CQ in both sub-bit levels. Figure 5 shows that AnTKV consistently outperforms CQ. Notably, when quantizing down to 0.375-bit, AnTKV maintains a tolerable performance degradation, with the average score dropping from 46.2 to 34.5.
@@ 4.4 Efficiency
In this experiment, we evaluate the efficiency of our AnTKV implementation against the FlashAttention-accelerated FP16 baseline on LLaMA-3-8B using a single A100-80GB GPU. As shown in Figure 6, AnTKV substantially extends the maximum context length from 128K to 384K. In long-context scenarios, we observe that intermediate activations consume a significant portion of memory. By applying a series of in-place operations, AnTKV ultimately supports up to 810K tokens under 1-bit quantization and 840K under 0.375-bit quantization. More importantly, AnTKV sustains low memory consumption while maintaining acceptable latency. Under 1-bit quantization, a noticeable increase in prefill latency is observed only at the 128K context length, which represents the upper limit supported by most current models. It is observed that 0.375-bit quantization introduces higher latency than its 1-bit. This increase in latency arises from the use of a larger number of centroids, which are employed to enhance model accuracy.
To evaluate decoding efficiency, we evaluate the throughput of AnTKV under a fixed context length of 1K tokens. As shown in Figure 7, AnTKV supports significantly larger batch sizes compared to the baseline method. By substantially reducing KV cache access, it improves throughput across all bit-widths. Particularly, the maximum throughput reaches 3.5x that of the baseline method under 1-bit quantization.
Figure 6: Memory and latency comparison. Bars show memory usage (red: 1-bit, green: 0.375-bit); lines indicate prefill latency across context lengths.
Figure 7: Decoding throughput comparison. Our method supports larger batch sizes and achieves higher throughput.
@@ 5 Limitation & Conclusion
Although AnTKV demonstrates its advantages in experiments, it also has several limitations. First, more accurate AnS for tokens and higher performance implementations for its computation may be possible. AnTKV demonstrates strong potential in LLM serving by substantially reducing the size of the KV cache, which in turn alleviates I/O and memory constraints to a significant extent. Nevertheless, further empirical validation is required.
We observed from several experiments that different tokens exhibit varying sensitivity to KV quantization, which opens a new opportunity for a better memory-accuracy trade-off. To this end, we present AnTKV, a two-stage token-aware VQ framework that involves offline gradient weighted centroid learning and online anchor token selection guided by AnS. Moreover, to ensure the efficiency of AnTKV in LLM inference, we implement a lightweight GPU kernel compatible with FlashAttention to compute AnS with minimal overhead. Extensive experiments show that AnTKV achieves superior performance under low and ultra-low-bit quantization, significantly extending the maximum context length (up to 840K tokens) on a single A100 GPU while maintaining low perplexity and high throughput. Our results demonstrate the potential of token-aware quantization in scaling LLM inference to longer contexts without sacrificing accuracy.
@@ Appendix A Distribution of pre- and post- RoPE Key
$\begin{split}\|\texttt{Attn}\left(\bm{Q},\bm{K}+\delta\bm{K},\bm{V}\right)-\texttt{Attn}\left(\bm{Q},\bm{K},\bm{V}\right)\|_{L_{1}}&\approx\|\left(\bm{I}_{n}\otimes\bm{V}^{T}\right)\texttt{Diag}\left(\texttt{Vec}(\bm{A})\right)\texttt{Vec}(\bm{Y})\|_{L_{1}}\\ &=\sum\limits_{i}\|\bm{V}^{T}\texttt{Diag}\left(\bm{A}_{i,:}\right)\left(\bm{I}_{n}-\bm{e}\bm{A}_{i,:}\right)\bm{X}_{i,:}^{T}\|_{L_{1}}\\ &\leq\sum\limits_{i}\sum\limits_{j}\|\left(\bm{V}^{T}\texttt{Diag}\left(\bm{A}_{i,:}\right)\left(\bm{I}_{n}-\bm{e}\bm{A}_{i,:}\right)\right)_{:,j}\|_{L_{1}}|\widetilde{\bm{Q}}_{i,:}\widetilde{\delta\bm{K}}_{j,:}^{T}|\\ &\leq\sum\limits_{j}\sum\limits_{i}\|\left(\bm{V}^{T}\texttt{Diag}\left(\bm{A}_{i,:}\right)\left(\bm{I}_{n}-\bm{e}\bm{A}_{i,:}\right)\right)_{:,j}\|_{L_{1}}\|{\bm{Q}}_{i,:}\|_{2}\|{\delta\bm{K}}_{j,:}\|_{L_{1}}.\\ \end{split}$ (9)
@@ Appendix E Computational Procedure of AnS in the Online Stage
[…]
@@ Appendix G Anchor Tokens Number Impact
To further investigate the impact of anchor token numbers on downstream tasks, we evaluated different anchor token numbers on the Trec and Qasper subtasks of LongBench under ultra-low-bit quantization settings. For convenience, we approximated $1\%$ of the anchor token number as 64. The results are shown in Figure 13. The results indicate that Trec’s performance increases linearly with the number of anchor tokens, eventually reaching a result within 10 points of the baseline when using $10\%$ anchor tokens. In contrast to Trec, Qasper’s accuracy remains stable with a small number of anchor tokens but shows a significant performance improvement when increased to $10\%$. We believe this difference stems from the inherent properties of each task.

### v2

@@ 3.2 Online Anchor Token Selection
In online inference, gradients can no longer serve as the token importance metric due to the cost imposed by real-time constraints. To address this, we perform an error propagation analysis of the attention operator (Attn). Specifically, the analysis derives a perturbation bound of Attn with respect to each row of $\bm{K}$ and $\bm{V}$, as presented in Theorem 3.1, with the detailed proof provided in Appendix B.
@@
We remark that the error propagation factors corresponding to $\bm{K}_{j,:}$ and $\bm{V}_{j,:}$ given in Theorem 3.1 can be regarded as the upper bound of the gradient of the attention operator related to $\bm{K}_{j,:}$ and $\bm{V}_{j,:}$.
The computation involving $\bm{K}$ in Theorem 3.1 introduces significant overhead and is therefore unsuitable for online inference. To address this limitation, we propose a simplified variant that excludes the contribution of $\bm{V}$ to the quantization error of $\bm{K}$. This leads to the following reformulation, with a detailed evaluation of AnS effectiveness provided in Appendix C.
$\text{AnS}(\bm{V}_{j,:})=\sum_{i}\bm{A}_{i,j}\quad\quad\text{AnS}(\bm{K}_{j,:})=\sum_{i}\bm{A}_{i,j}(1-\bm{A}_{i,j})\cdot\left\|\bm{Q}_{i,:}\right\|_{2}$ (5)
In online inference, during the prefill phase, AnS serves as an effective metric for identifying anchor tokens that induce substantial accuracy loss. In autoregressive decoding phase, AnS can still be computed. However, the anchor tokens it identifies may already have been quantized, which prevents preserving their full-precision values and limits error reduction. An important observation is that both AnS(K) and AnS(V) exhibit strong locality during the decoding phase (see Appendix D), with anchor tokens predominantly concentrated at the head and tail of the sequence. Experimental results show that the anchor tokens at the head of the sequence are consistently identified during the prefill stage, corresponding to sink tokens. This observation further demonstrates the effectiveness of our method. Building on the tail locality, we use a sliding-window approximation of AnS during decoding to further enhance efficiency while mitigating accuracy degradation.
@@ 3.3 Implementation
For the offline centroids learning stage, the gradients of $\bm{K}$ and $\bm{V}$ are employed as weights for centroid learning. We implement it with a custom LinearWithAct to capture KV cache and corresponding gradients. Subsequently, we employ the weighted k-means provided by cuML to perform efficient clustering.
In the online stage, AnS is derived from the error propagation factor given in Theorem 3.1 and Equation (5). To enable efficient long-context inference, we design and implement a dedicated GPU kernel using Triton that computes AnS in conjunction with FlashAttention. Because AnS requires reduction operations over the attention score matrix $\bm{A}$ and its transformed form $\bm{A}\odot(\bm{E}-\bm{A})$ along the query dimension (column-wise), direct fusion into FlashAttention is infeasible. To preserve the efficiency of FlashAttention, we decouple AnS computation and execute it immediately afterward. For this purpose, we extend FlashAttention to additionally output three auxiliary tensors: the $L_{2}$ norm of each query vector, the key-wise (row-wise) sum, and the key-wise maximum of the matrix $\bm{Q}\bm{K}^{T}$. These tensors allow the reconstruction of the attention scores and facilitate AnS computation with minimal overhead. Further implementation details are provided in Algorithm 1 of Appendix E.
Finally, since the application of RoPE disrupts the channel-wise magnitude distribution of $\bm{K}$ (see Appendix A), which otherwise exhibits large inter-cluster distances and small intra-cluster variances, the pre-RoPE strategy, consistent with hooper2024kvquant , is adopted in AnTKV.
@@ 4 Experiments
In this section, we present an extensive comparison between AnTKVand existing KV quantization methods. The experimental setup is detailed below.
Models, Datasets, Metrics, and Parameter Settings. To validate the effectiveness and generality of AnTKVin KV cache quantization, we evaluate five representative models from the LLaMA and Mistral families. For calibration, 128 samples of length 2048 are drawn from the WikiText2 training set. Model quality is assessed through three categories of benchmarks: (i) perplexity on WikiText-2 and C4; (ii) zero-shot accuracy on MMLU hendrycks2021mmlu , ARC-C clark2018arc , MathQA amini2019mathqa , and PIQA bisk2020piqa to evaluate understanding and reasoning; and (iii) long-context performance on LongBench bai2024longbench . For perplexity and zero-shot evaluations, quantized KV caches are directly used for attention outputs, whereas for LongBench, full precision KV cache is used to compute attention outputs and quantized KV cache is used during decoding. Across all benchmarks, anchor tokens are restricted to a small subset: 1% of the context length for perplexity, 16 for understanding and reasoning, and 64 for LongBench. For fair comparison, a sliding window of size 32 is applied in LongBench, following the mainstream setting.
Baselines. We compare AnTKVwith full precision and representative KV cache quantization methods, including KIVI liu2024kivi , SKVQ duanmu2024skvq , KVQuant-1% hooper2024kvquant , and CQ zhang2024kv . SKVQ is configured with a group size of 64 and five sink tokens xiaoefficient , while KVQuant retains four sink tokens. Since CQ results are not publicly available, we reproduced them following the methodology in their paper to the best of our understanding. For VQ settings, we adopt the notation “dnncmm”, covering 4-bit (d2m256), 2-bit (d4m256), 1-bit (d8m256), 0.75-bit (d16m4096), and 0.375-bit (d32m4096).
@@ 4.1 Perplexity Results
Perplexity is a standard benchmark that is widely used to evaluate the quality of the output of LLMs, with lower values indicating better performance. The perplexity results for different KV quantization approaches on WikiText-2 and C4 are presented in Table 1. The results in this table indicate that the proposed AnTKVconsistently achieves competitive or superior perplexity across various bit-widths and model architectures. Under 4-bit and 2-bit quantization, it achieves competitive performance compared to baseline. In the 1-bit and sub-bit regimes, it significantly outperforms all baselines. On the C4 dataset under sub-bit quantization, baseline methods suffer from extremely high perplexity, as fixed centroids fail to capture anchor tokens. By contrast, with its effective AnS design and anchor token selection, AnTKVsubstantially lowers perplexity, reducing it from 66.28 to 14.42 on LLaMA-3-8B at 0.75-bit.
Table 1: All evaluations are performed under the maximum context length of each model, specifically 4096 for LLaMA-2-7B and 8192 for LLaMA-3-8B and Mistral-7B. “Ours” refers to AnTKVwithout anchor tokens, whereas “Ours-1%” denotes AnTKVwith 1% of tokens designated as anchor tokens and retained in FP16. For clarity, the reported bit-widths exclude the contribution of centroids.
@@ 4.2 Understanding and Reasoning Benchmark
To assess the breadth of AnTKV’s understanding and reasoning capabilities, we evaluate it on four representative benchmarks using LLaMA-8B-Instruct and Mistral-7B-Instruct. These benchmarks target multi-domain knowledge reasoning (MMLU), complex question answering (ARC-Challenge), commonsense reasoning (PIQA), and mathematical problem solving (MathQA). Due to missing the implementations in the official repository, KIVI and KVQuant are not included. As shown, AnTKVconsistently maintains higher accuracy across LLaMA and Mistral models, with particularly strong advantages at 1-bit and sub-bit settings where baseline methods degrade sharply.
Figure 4: Evaluation of understanding and reasoning accuracy on MMLU, ARC-C, PIQA, and MathQA under different quantization bit-widths.
@@ 4.3 Long-Context Benchmark
To validate the effectiveness of AnTKV in handling long-context, We conduct several experiments on the LLaMA-8B-Instruct model using the LongBench benchmark, a diverse collection of tasks such as question answering, retrieval, and summarization, designed to systematically evaluate long-context understanding in language models. We report results on eleven representative sub-tasks from LongBench, along with averaged performance. Due to alignment issues of KIVI and SKVQ, we exclude the triviaqa and gov_report sub-tasks from the comparison. As shown in Figure 5, AnTKV preserves nearly FP16 at 4- and 2-bit quantization across almost all tasks. At the 1-bit quantization, the performance of KIVI and SKVQ has a significant drop. In contrast, AnTKV and CQ still maintain a relatively high accuracy. To further investigate the robustness under aggressive compression, we compare AnTKV and CQ in both sub-bit levels. Figure 5 shows that AnTKV consistently outperforms CQ. Notably, when quantizing down to 0.375-bit, AnTKV maintains a tolerable performance degradation, with the average score dropping from 46.2 to 38.1.
@@ 4.4 Efficiency
In this experiment, we evaluate the efficiency of our AnTKV implementation compared with huggingface baseline transformers on LLaMA-3-8B using a single A100-80GB GPU. As shown in Figure 6,  AnTKV substantially extends the maximum context length from 128K to 384K. In long-context inference, our profiling shows that intermediate activations account for a substantial portion of memory usage. By introducing a series of in-place operators, AnTKV supports up to 810K tokens under 1-bit quantization and 840K under 0.375-bit quantization, while maintaining low memory consumption. To evaluate decoding efficiency, we measure the throughput of AnTKV with a fixed context length of 1K tokens. As shown in Figure 7AnTKV enables substantially larger batch sizes and improves throughput across all bit-widths by reducing KV cache access. In particular, under 1-bit quantization, the maximum throughput reaches $3.5\times$.
Figure 6: KV cache memory size comparison. Gray bars denote full precision, red bars 1-bit, and green bars 0.375-bit quantization. Striped bars indicate results with in-place operators enabled.
Figure 7: Decoding throughput comparison. Our method supports larger batch sizes and achieves higher throughput. Dashed lines indicate results with in-place operators enabled.
@@ 5 Limitation & Conclusion
Although AnTKV demonstrates its advantages in experiments, it also has few limitations. First, more accurate AnS for tokens and higher performance implementations for its computation may be possible. AnTKV demonstrates strong potential in LLM serving by substantially reducing the size of the KV cache, which in turn alleviates I/O and memory constraints to a significant extent. Nevertheless, further empirical validation is required.
This work addresses the preservation of accuracy under ultra low bit KV cache quantization. We propose AnTKV , a vector quantization based framework that exploits intra vector correlations. AnTKV uses a dual stage design with offline token aware centroid learning and online anchor token selection, which mitigates the disproportionate error from anchor tokens. Across the LLaMA and Mistral families, AnTKV attains accuracy close to full precision and consistently surpasses baselines in the ultra low bit regime. It also scales LLaMA-3-8B to 840K tokens on a single 80 GB A100, and increases decoding throughput by up to $3.5\times$.
@@ Appendix A Distribution of pre- and post- RoPE Key
[…]
@@ Appendix E Computational Procedure of AnS in the Online Stage
8:   $\text{AnS}({\bm{K}})_{i}\leftarrow\text{AnS}({\bm{K}})_{i}+\texttt{col\_sum}(\bm{A}_{i,j}\cdot(1-\bm{A}_{i,j}),\ \text{row-wise})$
9:   $\text{AnS}({\bm{V}})_{i}\leftarrow\text{AnS}({\bm{V}})_{i}+\texttt{col\_sum}(\bm{A}_{i,j},\ \text{row-wise})$
@@ Appendix G Anchor Tokens Number Impact
To further investigate the impact of anchor token numbers on downstream tasks, we evaluated different anchor token numbers on the Trec and Qasper subtasks of LongBench under ultra-low-bit quantization settings. For convenience, we approximated $1\%$ of the anchor token number as 64. The results are shown in Figure 13. it illustrates that both the Trec and Qasper subtasks exhibit a consistent improvement pattern as the number of anchor tokens increases. In particular, moving from 0% to 1% anchor tokens leads to a substantial performance gain across all quantization settings, highlighting that even a very small proportion of anchor tokens can effectively mitigate the degradation introduced by ultra-low-bit quantization. Beyond this point, the improvements from 1% to 2%, 2% to 5%, and 5% to 10% follow an approximately linear trend, with performance gradually approaching the FP16 baseline. These results demonstrate that anchor tokens play a dual role that a small fraction is sufficient to deliver immediate and significant benefits, while larger allocations further provide steady, near-linear enhancements in downstream task performance.